OpenStudy (anonymous):

How many integer values are there for k for which 4x^2+kxy-9y^2 is factorable?

OpenStudy (dumbcow):

i say 9

OpenStudy (anonymous):

I did this problem a while back. hold on it's in my posts

OpenStudy (anonymous):

How do you get 9?

OpenStudy (mathstudent55):

Use the ac method and factor by grouping. ac = 4 * (-9) = -36 -36 = -1 * 36 -1 + 36 = 35 -36 = -36 * 1 -36 + 1 = -35 -36 = -2 * 18 -2 + 18 = 16 -36 = -18 * 2 -18 + 2 = -16 -36 = -3 * 12 -3 + 12 = 9 -36 = -12 * 3 -12 + 3 = -9 -36 = -4 * 9 -4 + 9 = 5 -36 = -9 * 4 -9 + 4 = -5 -36 = -6 * 6 -6 + 6 = 0 Each of the sums above in the right column is a value of k that will make this factorable.

OpenStudy (dumbcow):

i was going to let @mathstudent55 explain that :)

OpenStudy (anonymous):

ohh i understand now, thanks!

OpenStudy (mathstudent55):

wlcm